OpenStudy (anonymous):

"The differential equation is not a polynomial equation in its derivatives and so its degree is not defined." What does "polynomial equation in its derivatives" mean? I don't understand, how do we decide when the degree of a differential equation is undefined.

OpenStudy (amistre64):

a polynomial degree is defined by its highest exponent. The degree of a differential equation is not related to the degrees of its polynomial parts.

OpenStudy (anonymous):

There are certain cases when the degree of a differential equation is undefined. How do I determine that in a differential equation the degree is undefined?

OpenStudy (amistre64):

hmm, id have to google that one ... i do know that: order is the higest derivative; and degree is the exponent of the highest order [y^(4)]^2 + y1 = sin(x) has order 4, and degree 2; but when is degree undefined? maybe when it is a variable?

OpenStudy (amistre64):

yeah, it appears that if something other than a numerical exponent is playing with a derivative part; then the degree cannot be determined

OpenStudy (anonymous):

@amistre64 I have figured out that the degree is undefined when either the exponent is a derivative or there is something like sin(dy/dx) given. Just as x^2 + 1 = 0 is a polynomial but x^2 + sin(x) =0 is not

OpenStudy (amistre64):

sounds reasonable to me :)